After 5 seconds, what is the initial velocity of a body thrown upward, if, above its starting place, it is to be 50 𝑓𝑑 high?


1
Expert's answer
2021-09-24T09:27:03-0400

Find the initial velocity at which the body was thrown:


"h=v_0t-\frac{gt^2}{2},\\\space\\
v_0=\frac{h}{t}+\frac{gt}{2}=90\text{ ft/s}."